# The exact travelling wave solutions of a KPP equation

### v5

@@ Abstract
We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed. Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
@@ I Introduction
where $\kappa(>0)$ is the carrying capacity of the environment. The term $f(u)=ku(1-u/\kappa)$ represents a nonlinear growth rate which is proportional to $u$ for small $u$, but decreases as $u$ increases, and vanishes when $u=\kappa$. Equation (1) arises in many physical, biological, and chemical problems involving diffusion and nonlinear growth. It is convenient to introduce the nondimensional quantities $x^{*},t^{*},u^{*}$, so after dropping the asterisks, equation (1) takes the nondimensional form
@@
where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=F(u)$ (5)
@@
Equation (5) is a nonlinear nonintegrable differential equation. The Cauchy problem cannot be solved by the inverse scattering transform for this equation and the problem of finding some exact solutions for this equation is an important task. Several methods of obtaining such solutions has been developped. We will mention here a few of them: the tanh-function method 4 ; 5 ; 6 , the Jacobi elliptic function expansion method 8 , the simplest equation method 11 ; 12 , the Exp-function method 14 , the $G^{\prime}/G$ expansion method 15 ; 16 ; 17 ; 18 , and the Kudryashov method 19 .
In this paper we will use the method of exclusion of the independent variable benguria ; kogan3 ; kogan4 ; kogan5 , which allows to reduce the problem of integration of the second order ODE to the problem of integration of the first order ODE, to obtain the exact solutions of Eq. (5). Thus we were able to obtain the exact solutions of the KPP equation which to the best of our knowledge were not known before. Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
@@ II The ODE which doesn’t contain explicitly the independent variable
where $n>0$ and $a\geq 0$ are some constants, so (5) takes the form
$u_{\xi\xi}+cu_{\xi}=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (12)
Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (12) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
Starting from (12), we obtain Eq. (9) in the form
$pp_{u}+cp=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (13)
One can easily check up that there exists the solution of (13) satisfying the boundary conditions (10)
@@
provided that the speed of the wave is
@@
Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (12) as
@@
(An alternative way to obtain (17) was presented in our previous publication kogan4 .)
@@
In Appendix A we’ll see how the results (14) - (17) can be generalized in the case when advection is taken into account in addition to diffusion. In Appendix C we’ll analyze the solution of (13) (for $n=1$) for $c$ close to $c_{0}$.
@@
In the previous Section we postulated two boundary conditions (10) for (9). However, the equation is of the first order. The first impression is that the problem is overdetermined, and the solution obtained there exists due to the finetuning of the l.h.s. of the equation. So one may think that we should finetune the speed $c$ for the solution of (13) to exist, and the solution will not exist for a general value of $c$. In fact, didn’t we finetune $c$ in the previous Section? However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (13) for arbitrary $c\gg 1$ is presented. It looks like we have a problem.
To solve this problem we have to return to Eq. (5) with the boundary conditions (6). Actually, the mechanical analogy allows us easily visualize the solution of the equation, but let us forget for the time being about this analogy and look at the equation purely mathematically. We immediately realize that the boundary conditions (6) - they are exceptional goursat $u_{\xi\xi}$ is not defined at $\xi=\pm\infty$. To make use of the boundary conditions consider Eq. (5) for $\xi\to\pm\infty$. For $\xi\to+\infty$ the equation can be presented as
$u_{\xi\xi}+cu_{\xi\xi}+F^{\prime}(0)u=0.$ (19)
For $\xi\to-\infty$ the equation can be presented as
@@
After we understood the true nature of the boundary conditions (6) for (5), the situation with the boundary conditions (10) for Eq. (9) becomes obvious. Note that the boundary conditions (10) are also exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
In the vicinity of $u=1$, Eq. (9) can be simplified to
@@ Acknowledgements.
We are grateful to G. G. Contreras, A. Gupta, M. Ma, B. Malomed, C. Ou and R. Selvaraj for the insightful discussions.
Appendix A Generalized Burgers–Huxley equation
@@
Appendix B Perturbation theory I
@@
Appendix C Perturbation theory II
In Section II we have solved Eq. (13) for $c=c_{0}$. Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series. To make life simpler $n=1$, we’ll consider only the case $n=1$. Presenting the solution as

### v6

@@ Abstract
We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed. Additionally we obtain the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
@@ I Introduction
where $\kappa(>0)$ is the carrying capacity of the environment. The term $f(u)=ku(1-u/\kappa)$ represents a nonlinear growth rate which is proportional to $u$ for small $u$, but decreases as $u$ increases, and vanishes when $u=\kappa$. Equation (1) arises in many physical, biological, and chemical problems involving diffusion and nonlinear growth. It is convenient to introduce the nondimensional quantities
$x^{*}$ $=x\sqrt{k/\nu},$ (2a)
$t^{*}$ $=kt,$ (2b)
$u^{*}$ $=u/\kappa,$ (2c)
so after dropping the asterisks, equation (1) takes the nondimensional form
@@
where $c$ is the wave speed (measured in the units of $\sqrt{\nu/k}$), and $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=-F(u)$ (6)
@@
Equation (6) is a nonlinear nonintegrable differential equation. The Cauchy problem cannot be solved by the inverse scattering transform for this equation and the problem of finding some exact solutions for this equation is an important task. Several methods of obtaining such solutions has been developed. We will mention here a few of them: the tanh-function method 4 ; 5 ; 6 , the Jacobi elliptic function expansion method 8 , the simplest equation method 11 ; 12 , the Exp-function method 14 , the $G^{\prime}/G$ expansion method 15 ; 16 ; 17 ; 18 , and the Kudryashov method 19 .
In this paper we will use the method of exclusion of the independent variable benguria ; rosu ; kogan3 ; kogan4 ; kogan5 , which allows to reduce the problem of integration of the second order ODE to the problem of integration of the first order ODE, to obtain the exact solutions of Eq. (6). Thus we were able to obtain the exact solutions of the KPP equation which to the best of our knowledge were not known before. Additionally we obtain the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
@@ II The ODE which doesn’t contain explicitly the independent variable
where $n>0$ and $a\geq 0$ are some constants, so (6) will take the form
$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (13)
Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (13) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath . Starting from (13), we obtain Eq. (10) in the form
$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (14)
In search of elementary solutions of (14) let us return to Eq. (10), and in accordance with the well-known in mathematics principle, stating that the more general the problem is, the easier it is to solve it, let us even generalize it to
$pp_{u}+\gamma(u)p=-F(u),$ (15)
where $\gamma(u)$ is some function. Being inspired by the method of factorization gonzales , we present the r.h.s. of (15) as
$-F(u)=\Phi(u)\Psi(u).$ (16)
After that we realize that there exists the particular solution of (15)
$p(u)=P\Phi(u),$ (17)
provided the unknown multiplier $P$ and the function $\gamma(u)$ satisfy equation
$P^{2}\Phi_{u}+P\gamma(p)=\Psi(u).$ (18)
This includes the case of nonlinear $F(u)$ considered in Ref. an
$F(U)=F^{\prime}(0)\left[u-A(u)\right]\left[1+\chi A^{\prime}(u)\right],$ (19)
where $A(u)$ is some function, $\chi$ is some numerical parameter, and $\gamma(p)=c=$ const. In fact, taking
$\Phi(u)=F^{\prime}(0)\left[A(u)-u\right]$ (20)
we obtain (18) in the form
$P^{2}F^{\prime}(0)\left[A^{\prime}(u)-1\right]+Pc=1+\chi A^{\prime}(u).$ (21)
We see that choosing $P$ and $c$ we can turn (21) into the identity.
Now let us return to Eq. (14) and take
$\Phi(u)=u\left(u^{n}-1\right),\hskip 28.45274pt\Psi(u)=\left(u^{n}+a\right).$ (22)
Equation (18) takes the form
$P^{2}\left[(n+1)u^{n}-1\right]+Pc=u^{n}+a.$ (23)
We immediately realize that the solution of (14) satisfying the boundary conditions (11)
@@
exists if the speed of the wave is
@@
Substituting thus obtained $p(u)$ into (9) and integrating we obtain the solution of (13)
@@
To additionally illustrate the method of integration used in this Section, in the Appendix A we obtain in the framework of the method the exact solution of the generalized Duffing-Van der Pol equation.
In Appendix B we’ll see how the results (24) - (27) can be generalized in the case when advection is taken into account in addition to diffusion.
The connection between the existence of the elementary solution of (10) and the possible symmetry of the equation will be analysed in our next publication.
@@
In the previous Section we postulated two boundary conditions (11) for (10). However, the equation is of the first order. The first impression is that the problem is overdetermined, and the existence of the solution of (10) demands the finetuning of the l.h.s. of the equation. Being more specific, one may think that the solution of (14) will not exist for a general value of $c$. However, the last statement immediately comes into contradiction with the results of the Appendix C, where the solution of (14) for arbitrary $c\gg 1$ is presented. It looks like we have a problem.
To solve this problem we have to return to Eq. (6) with the boundary conditions (7). The boundary conditions are exceptional goursat $u_{\xi\xi}$ is not defined at $\xi=\pm\infty$. However, for $\xi\to+\infty$ the equation asymptotically becomes
$u_{\xi\xi}+cu_{\xi\xi}+F^{\prime}(0)u=0,$ (29)
and for $\xi\to-\infty$
@@
After we understood the true nature of the boundary conditions (7) for (6), the situation with the boundary conditions (11) for Eq. (10) - which are also exceptional - becomes obvious.
In the vicinity of $u=1$, Eq. (10) becomes asimptotically
@@ Acknowledgements.
We are grateful to G. Gonzales, A. Gupta, M. Ma, B. Malomed, A. Mogilner, H. C. Rosu, C. Ou and R. Selvaraj for the insightful discussions.
Appendix A Exact solution of the generalized Duffing –Van der Pol equation
Consider the generalized Duffing –Van der Pol equation
$X_{tt}+\left(a-bF_{X}\right)X_{t}=F(X),$ (45)
where $X(t)$ is a deviation from the balance of the oscillator, $t$ is time, $F(X)$ is the force field determined by the model, and $a,b$ are the parameters determining the generalized friction. (Equation (45) describes, in particular, travelling waves in a lossy transmission line with nonlinear capacitance kogan3 .) Introducing new dependent variable
$p=X_{t}$ (46)
and considering $X$ as the independent variable, we reduce (45) to the first order ODE
$pp_{X}+\left(a-bF_{X}\right)p=F(X).$ (47)
The particular solution of (47)
$p(X)=bF(X)$ (48)
exists if
$ab=1.$ (49)
Looking back at Section II we realize that the solution corresponds to the factorization
$F(X)=F(X)\cdot 1.$ (50)
Substituting thus obtained function $p(X)$ into (46) we obtain the solution of (45)
$t=\int\frac{dX}{bF(X)}.$ (51)
Consider $F(X)$ which has two adjacent zeros at $X_{1}$ and $X_{2}$ ($F(X_{1})=F(X_{2})=0$), one of these points being the point of maximum of the potential $V(X)$ defined by the force field $F(X)$, the other - the point of minimum. Then (51) describes explicitly the motion of the particle in the potential well $V(X)$ which starts in the infinite past at the potential maximum and ends in the infinite future at the potential minimum. Finetuning of the friction (Eq. (49)) is necessary for the absence of oscillations in the vicinity of the potential minimum, which gives the opportunity to integrate (45) in quadratures.
Let us be even more specific and consider the Duffing –Van der Pol equation from Ref. 19
$X_{tt}+\left(\alpha+\beta X^{2}\right)X_{t}=gX-X^{3}.$ (52)
Equation (52) is the particular case of (45) corresponding to
$F(X)$ $=gX-X^{3},$ (53a)
$a$ $=\alpha+\beta g/3$ (53b)
$b$ $=\beta/3.$ (53c)
Thus the condition (49) takes the form
$\alpha=3/\beta-\beta g/3;$ (54)
integration in (51) can be easily performed and we reproduce the result of Ref. 19
$X(t)=\frac{\sqrt{g}}{\sqrt{1-e^{-2\beta gt/3}}}.$ (55)
Appendix B Generalized Burgers–Huxley equation
@@
Appendix C Perturbation theory I
@@
Appendix D Perturbation theory II
In Section II we have solved Eq. (13) for $c=c_{0}$. Motivated by the perturbation theory of the previous Section, let us study the solution of (13) for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series. To make life simpler $n=1$, we’ll consider only the case $n=1$. Presenting the solution of (14) (for $n=1$) as